1. The height of a triangle is 4 feet less than the base. The area of the triangle is 16 square feet. Find the length of the base and height of the triangle.
2. The perimeter of a rectangle is 32 inches. The area of the rectangle is 48 square inches. Find the dimensions of the rectangle.
3. Find the radius of a cylinder whose volume is 2512 cubic centimeters and whose height is 8 centimeters.




Answer by checkley77(12569)

the formula is bh/2=area
b*(b-4)/2=16
(b^2-4b)/2=16
b^2-4b=32
b^2-4b-32=0
(b-8)(b+4)=0
b-8=0
b=8 feet for the base.
h=8-4=4 for the height.
proof
8*4/2=16
32/2=16
16=16